OpenStudy (anonymous):

In a high school that enrolls only tenth, eleven, and 12 graders, there are 1,134, tenth graders, 1,052 eleven graders and 983 twelfth graders enrolled in the school. A). On a particular day 8% of the student are absent. What is the approximate total number of student that are present for the day?

OpenStudy (radar):

Find total number of students when all are present. 1,134+1,052+983=3,169 If 8% are absent, then we can assume that 92% are present. Find out what 92% of 3,169 students are

OpenStudy (radar):

92% is 92/100 or .92

OpenStudy (radar):

Solve .92 x 3,169

OpenStudy (anonymous):

Is that correct?

OpenStudy (anonymous):

2915.48

OpenStudy (radar):

Yes, but obviously we have to round that off to the nearest whole number, can't have a portion of a student present. I would go with the 2,195 student.

OpenStudy (anonymous):

Ok, Can you help me with part B of that question please.

OpenStudy (anonymous):

On the same day 60% of the students who were present chose to buy lunch in the cafeteria. Suppose 20% of the students who bought their lunch at the cafeteria bough hot lunch meals, while the other students who bought lunch in the cafeteria that day purchased salads or fast foods. What percent of the students who were present that day purchased salads or fast foods?

OpenStudy (radar):

I meant to say 2,915 students

OpenStudy (anonymous):

i got it

OpenStudy (radar):

OK continuing on for that day where there was 2915 students. 60% of those students ate in the cafeteria, or .60 X 2915 = 1749. 20% of this number bought the hot lunch. this means 1749 x .20 =350 had hot lunches. That means 1749 -350 = 1399 had salad. The question is what percent of students attending that day had salads/fast food. We have already computed the total students present that day as 2195 students. The task at hand is to determine what percent of 2915 total students is the 1399 salad students. This is set up like a fraction with the 1399 (the portion) as numerator and the 2195 (the whole) as the denominator or 1399/2915=.4799 rounding off to .48 convert to percent by moving the decimal two places to the right (multiplying by 100) and add the percent sign: 48%

OpenStudy (anonymous):

Hi. Are you there still?

OpenStudy (anonymous):

Radar

OpenStudy (radar):

I am back but have to leave. Good luck with your problems